OpenStudy (anonymous):

5 + 3 = x^3 – y^3 x^2 +xy + y^2

OpenStudy (anonymous):

Is this problem\[\frac{5}{x^3-y^3}+\frac{3}{x^2+xy+y^2}=\]

OpenStudy (anonymous):

If it is,\[\frac{5}{x^3 – y^3}+\frac{3}{x^2 +xy + y^2}=\frac{5}{(x-y)(x^2 +xy + y^2)}+\frac{3}{x^2 +xy + y^2}\]\[=\frac{5+3x-3y}{x^3-y^3}\]

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

No sweat. Do math every day.